OpenStudy (anonymous):

HELP PLEASE! ???? Auditions for a new, upcoming band in Chicago were held with 72 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 34, how many women auditioned? 22 44 50 60

OpenStudy (anonymous):

x + y = 72 2x -3y = 34

OpenStudy (anonymous):

50 women

OpenStudy (anonymous):

Okay, thanks!!!! :)

OpenStudy (anonymous):

y = women x = men

OpenStudy (anonymous):

its actually 22 women

OpenStudy (anonymous):

M + W = 72 2M - 3W = 34 Using the first equation: M=72-W Now substitute this into the second equation: 2M-3W=34 2(72-W)-3W=34 144-2W-3W=34 -5W=-110 Divide both sides by -5 and you get: W=22 (by the way W=Women and M=Male)